### v1

@@ 4.2 Bounds Derived from Linear Programming
To incorporate many complicated settings, suppose that $W(d)=A_{d}q$ and $p=Bq$ for some known row vector $A_{d}$ and matrix $B$, an unknown vector $q$ in a simplex $\mathcal{Q}$, and a vector $p$ that is estimable from data. Typically $q$ is a vector of probabilities of a latent variable that governs the DGP; see Balke and Pearl, (1997, 2011), Mogstad et al., (2018), Han, (2024) and Han and Yang, (2024). The linearity in this assumption is usually implied by the nature of a particular problem (e.g., discreteness). Then we have
@@
An example of this setting is as follows. Let $(Y,D,Z)$ be a vector of a binary outcome, treatment and instrument and let $p$ be a vector with entries $p(y,d|z)\equiv P(Y=y,D=d|Z=z)$ across $(y,d,z)\in\{0,1\}^{3}$. [^5] See Han and Yang (2024) for the use of linear programming with continuous $Y$. Suppose $W(d)=E[Y(d)]$ for $d\in\{0,1\}$. Then, we can define the response type $\varepsilon\equiv(Y(1),Y(0),D(1),D(0))$ with a realized value $e\equiv(y(1),y(0),d(1),d(0))$, where $Y(d)$ denotes the potential outcome under treatment $d$ and $D(z)$ denotes the potential treatment under instrument value $z$. Let $q(e)\equiv P(\varepsilon=e)$ be the latent distribution. Then
$W(d)$ $=\Pr[Y(d)=1]=\sum_{e:y(d)=1}q(e)\equiv A_{d}q,$
@@
where $\tilde{B}\equiv\left[\begin{array}[]{c}B\\ \boldsymbol{1}^{\prime}\end{array}\right]$ is a $(d_{p}+1)\times d_{q}$ matrix with $\boldsymbol{1}$ being a $d_{q}\times 1$ vector of ones, and $\tilde{p}\equiv\left[\begin{array}[]{c}p\\ 1\end{array}\right]$ is a $(d_{p}+1)\times 1$ vector. By using a vertex enumeration algorithm (e.g., Avis and Fukuda (1992)), one can find all (or a relevant subset) of vertices of the polyhedra $\{\lambda:\tilde{B}^{\prime}\lambda\geq-A_{d}^{\prime}\}$ and $\{\lambda:\tilde{B}^{\prime}\lambda\geq A_{d}^{\prime}\}$. Let $\Lambda_{L,d}\equiv\{\lambda_{1},...,\lambda_{J_{L,d}}\}$ and $\Lambda_{U,d}\equiv\{\lambda_{1},...,\lambda_{J_{U,d}}\}$ be the sets that collect such vertices, respectively. Then, it is easy to see that $L(d)=\max_{\lambda\in\Lambda_{L,d}}-\tilde{p}^{\prime}\lambda$ and $U(d)=\min_{\lambda\in\Lambda_{U,d}}\tilde{p}^{\prime}\lambda$, and thus Assumption 3.1 holds.
@@
where $\Delta_{\tilde{d},{d}}\equiv A_{\tilde{d}}-A_{{d}}$. Analogous to the vertex enumeration argument above, let $\Lambda_{L,\tilde{d},{d}}\equiv\{\lambda_{1},...,\lambda_{J_{L,\tilde{d},{d}}}\}$ and $\Lambda_{U,\tilde{d},{d}}\equiv\{\lambda_{1},...,\lambda_{J_{U,\tilde{d},{d}}}\}$ be the sets that collect all (or a relevant subset) of vertices of the polyhedra $\{\lambda:\tilde{B}^{\prime}\lambda\geq-\Delta_{\tilde{d},{d}}^{\prime}\}$ and $\{\lambda:\tilde{B}^{\prime}\lambda\geq\Delta_{\tilde{d},{d}}^{\prime}\}$, respectively. Then, $L(\tilde{d},{d})=\max_{\lambda\in\Lambda_{L,\tilde{d},{d}}}-\tilde{p}^{\prime}\lambda$ and $U(\tilde{d},{d})=\min_{\lambda\in\Lambda_{U,\tilde{d},{d}}}\tilde{p}^{\prime}\lambda$. Let $\widehat{\mathcal{D}}=\{d:\widehat{L}(\tilde{d},{d})\leq 0,\forall\tilde{d}\neq d\}$ where $\widehat{L}(\tilde{d},{d})$ is the sample counterpart of $L(\tilde{d},{d})$ with $\widehat{\tilde{p}}\equiv\left[\begin{array}[]{c}\hat{p}\\ 1\end{array}\right]$ replacing $\tilde{p}\equiv\left[\begin{array}[]{c}p\\ 1\end{array}\right]$. Partition $\lambda$ as $\lambda=(\lambda^{1\prime},\lambda^{0})^{\prime}$ where $\lambda^{0}$ is the last element of $\lambda$. Note that $d\in\widehat{\mathcal{D}}$ if and only if
@@ 4.3 Empirical Welfare Maximization with Observational Data
Consider allocating a binary treatment based on observed covariates $X\in\mathcal{X}$. A treatment allocation rule can be defined as a function $\delta:\mathcal{X}\rightarrow\{0,1\}$ in a class of rules $\mathcal{D}$. Consider the utilitarian welfare of deploying $\delta$ relative to treating no one (Manski (2009)): $W(\delta)=E[Y(\delta(X))-Y(0)]$. The optimal allocation $\delta^{*}$ satisfies
@@
Note that
$E[Y(\delta(X))-Y(0)]$ $=E\left[\delta(X)\Delta(X)\right],$
where $\Delta(X)\equiv E[Y(1)-Y(0)|X]$. This problem is considered in Kitagawa and Tetenov, (2018) and Athey and Wager (2021), among others. When only observational data for $(Y,D,X)$ are available with $D$ being endogenous, $W(\delta)$ is only partially identified unless strong treatment effect homogeneity is assumed. This problem has been studied in Kallus and Zhou (2021), Pu and Zhang (2021), D’Adamo (2021) and Byambadalai (2021), among others. Using instrumental variables, one can consider bounds on the conditional ATE based on conditional versions of the bounds considered in Sections 4.1 and 4.2 (i.e., Manski’s bounds or bounds produced by linear programming).
@@
$L(\delta)$ $\equiv E[\delta(X)L(X)],$
$U(\delta)$ $\equiv E[\delta(X)U(X)].$
Similarly, bounds on $W(\tilde{\delta})-W({\delta})=E[(\tilde{\delta}(X)-{\delta}(X))\Delta(X)]$ can be characterized as
$L(\tilde{\delta},{\delta})$ $\equiv E[(\tilde{\delta}(X)-{\delta}(X))L(X)],$ (4.7)
$U(\tilde{\delta},{\delta})$ $\equiv E[(\tilde{\delta}(X)-{\delta}(X))U(X)].$
@@
Suppose $\mathcal{X}$ is finite and $\mathcal{X}=\{x_{1},...,x_{K}\}$ where $x_{k}$ can be a vector and $K$ can potentially be large. For simplicity of exposition, suppose $\mathcal{X}=\{0,1,2\}$. Then $\mathcal{D}=\{\delta_{1},...,\delta_{8}\}$ where each $\delta_{j}$ corresponds to a mapping type from $\{0,1,2\}$ to $\{0,1\}$. To verify Assumptions 3.1 and 3.2, we proceed as follows. For given $x\in\mathcal{X}$, by arguments analogous to those in Sections 4.1 and 4.2, bounds $L_{x}$ and $U_{x}$ on $\Delta(x)$ satisfy, for some scalars ~$\tilde{\ell}_{j}$ and $\tilde{u}_{j}$ and row vectors $\ell_{j}$ and $u_{j}$,
$L_{x}$ $=\max_{j\in\{1,...,J_{L}\}}\{\tilde{\ell}_{j}+\ell_{j}p(x)\},$
$U_{x}$ $=\min_{j\in\{1,...,J_{U}\}}\{\tilde{u}_{j}+u_{j}p(x)\},$
where $p(x)$ is the vector of $p(y,d|z,x)$’s across $(y,d,z)$. Then, by Jensen’s inequality, for each $\delta\in\mathcal{D}$,
$L(\delta)$ $\geq\tilde{L}(\delta)\equiv\max_{j\in\{1,...,J_{L}\}}\{\tilde{\ell}_{j}E[\delta(X)]+\ell_{j}E[\delta(X)p(X)]\},$
$U(\delta)$ $\leq\tilde{U}(\delta)\equiv\min_{j\in\{1,...,J_{U}\}}\{\tilde{u}_{j}E[\delta(X)]+u_{j}E[\delta(X)p(X)]\}.$
Note that $\tilde{L}(\delta)$ and $\tilde{U}(\delta)$ are non-sharp bounds; for calculation of sharp bounds, see Section 4.4. We can verify Assumption 3.1 with $\tilde{L}(\delta)$ and $\tilde{U}(\delta)$ by defining
$p$ $=\left(\begin{array}[]{c}E[\delta_{1}(X)]\\ E[\delta_{1}(X)p(X)]\\ \vdots\\ E[\delta_{8}(X)]\\ E[\delta_{8}(X)p(X)]\end{array}\right)$
and, for $\delta=\delta_{1}$ as an example, by using
$\ell_{\delta_{1},j}$ =(~$=(\begin{array}[]{ccccc}\tilde{\ell}_{j}&\ell_{j}&0&\cdots&0\end{array}).$
Similarly, we can verify Assumptions 3.2 and 3.3 by estimating $p(X)$ and $E[\delta(X)]$ with sample means and ${E}[\delta(X){p}(X)]$ with $\frac{1}{n}\sum_{i}^{n}\delta(X_{i})\hat{p}(X_{i}).$ If the data $\{Y_{i},D_{i},Z_{i},X_{i}\}$${}_{i=1}^{n}$ form a random sample, and $\widehat{\mathcal{D}}=\{\delta\in\mathcal{D}:\widehat{L}(\tilde{\delta},\delta)\leq 0\,\forall\tilde{\delta}\neq\delta\}$ for $\widehat{L}(\tilde{\delta},\delta)$ defined the same as ${L}(\tilde{\delta},\delta)$ in (4.7) after substituting $\hat{p}$ for $p$, Assumptions 3.2 and 3.3 hold.
@@
4.4 Empirical Welfare Maximization via Linear Programming
Continuing from Section 4.3, we show how sharp bounds on $W(\delta)$ can be computed using linear programming. This can be done by extending the example in Section 4.2 with binary $Y$. Again, let $\mathcal{X}=\{x_{1},...,x_{K}\}$. Let $q(e|x)\equiv P(\varepsilon=e|X=x)$. In analogy,
$E[Y(d)|X=x]$ $=\Pr[Y(d)=1|X=x]=\sum_{e:y(d)=1}q(e|x)\equiv A_{d}q(x),$
where $q(x)$ is a vector with entries $q(e|x)$ across $e$, and
$\Pr[Y=1,D=d|Z=z,X=x]=\Pr[Y(d)=1,D(z)=d|X=x]$
$=\sum_{e:y(d)=1,d(z)=d}\Pr[\varepsilon=e|X=x]\equiv B_{d,z}q(x),$
$Bq(x)$ $=p(x).$
Now we can construct a linear program for welfare:
$W(\delta)=E[\delta(X)\Delta(X)]$ $=\sum_{x_{k}\in\mathcal{X}}p(x_{k})\delta(x_{k})\Delta(x_{k})$
$=\sum_{x_{k}\in\mathcal{X}}p(x_{k})\delta(x_{k})(A_{1}-A_{0})q(x_{k}).$
Therefore $W(\delta)$ satisfies the structure of Section 4.2 and by analogous arguments, Assumptions 3.13.3 hold.
@@ 4.5 Bounds for Dynamic Treatment Effects
Let $T=2$. Then $Y\equiv(Y_{1},Y_{2})$ and $D\equiv(D_{1},D_{2})$. We are interested in the optimal policy $d^{*}$ that satisfies
@@
where $\mathcal{D}\equiv\{(1,1),(1,0),(0,1),(0,0)\}$. The sign of the welfare difference, $W(d)-W(\tilde{d})$ for $d,\tilde{d}\in\mathcal{D}$, is useful for establishing the ordering of $W(d)$ with respect to $d$ and thus to identifying $d^{*}$. However, without additional identifying assumptions, we can only establish a partial ordering of $W(d)$ based on the bounds on the welfare difference. This will produce the identified set $\mathcal{D}^{*}$ for $d^{*}$.
An example of the welfare is $W(d)\equiv E[Y_{2}(d)]$, namely, the average potential terminal outcome. The bounds on welfare $W(d)$ are
$L(d)$ $\equiv\max_{z}L(d;z),$ (4.8)
$U(d)$ $\equiv\min_{z}U(d;z),$ (4.9)
@@
$L(d;z)$ $\equiv P(Y_{2}=1,D=d|Z=z),$
$U(d;z)$ $\equiv P(Y_{2}=1,D=d|Z=z)+\sum_{d^{\prime}\neq d}P(D=d^{\prime}|Z=z),$
@@
$W(\tilde{d})-W({d})=E[Y_{2}(\tilde{d})-Y_{2}({d})]$ $=P(Y_{2}(\tilde{d})=1)-P(Y_{2}({d})=1).$
@@
$L(\tilde{d},{d})$ $\equiv L(\tilde{d})-U({d}),$ (4.10)
$U(\tilde{d},{d})$ $\equiv U(\tilde{d})-L({d}).$ (4.11)
Another example of welfare is the joint distribution $W(d)\equiv P(Y(d)=(1,1))$ where $Y(d)\equiv(Y_{1}(d_{1}),Y_{2}(d))$. The bounds on $W(d)$ in this case are $L(d)\equiv\max_{z}L(d;z)$ and $U(d)\equiv\min_{z}U(d;z)$ where
$L(d;z)$ $\equiv P(Y_{2}=1,D=d|Z=z),$
$U(d;z)$ $\equiv P(Y_{2}=1,D=d|Z=z)+P(Y_{1}=1,D_{1}=d_{1},D_{2}=d_{2}^{\prime}|Z=z)$
$\quad+P(D_{1}=d_{1}^{\prime},D_{2}=d_{2}|Z=z)+P(D_{1}=d_{1}^{\prime},D_{2}=d_{2}^{\prime}|Z=z).$
Consider the effect of treatment on the joint distribution, for example,
$W(1,1)-W(1,0)$ $=P(Y(1,1)=(1,1))-P(Y(1,0)=(1,1)).$
@@
$L(\tilde{d},{d})$ $\equiv\max_{z}P(Y_{2}=1,D=(1,1)|Z=z)$
$\quad-\min_{z}[P(Y_{2}=1,D=(1,0)|Z=z)+P(Y_{1}=1,D=(1,1)|Z=z)$
$\quad+P(D=(0,1)|Z=z)+P(D=(0,0)|Z=z)],$
$U(\tilde{d},{d})$ $\equiv\min_{z}[P(Y_{2}=1,D=(1,1)|Z=z)+P(Y_{1}=1,D=(1,0)|Z=z)$
$\quad+P(D=(0,1)|Z=z)+P(D=(0,0)|Z=z)]$
$\quad-\max_{z}P(Y_{2}=1,D=(1,0)|Z=z).$
@@
$\mathcal{D}^{*}$ =$=\{d:\nexists\tilde{d}\neq d\text{ such that }L(\tilde{d},d)>0\}$
=$=\{d:L(\tilde{d},d)\leq 0,\forall\tilde{d}\neq d\}.$
These examples are special cases of the model in Han, (2024). [^6] See also Han and Lee (2024) for other examples of dynamic causal parameters that can be used to define optimal treatments.
In both cases, it is easy to see that $L(d)$ and $U(d)$ satisfy Assumption 3.1 with $p$ being the vector of probabilities $p(y,d|z)\equiv P(Y=y,D=d|Z=z)$. To verify Assumption 3.2, let $\widehat{L}(\tilde{d},d)$ be the estimator of $L(\tilde{d},d)$ where the sample frequency replaces the population probability. Then $\widehat{\mathcal{D}}=\{d:\widehat{L}(\tilde{d},d)\leq 0,\forall\tilde{d}\neq d\}$.
@@
[…]
@@
and, for example, $\widehat{z}(d)=1$ if and only if
$\widehat{L}(d;0)-\widehat{L}(d;1)$ $\leq 0.$
A similar formulation follows for the second example. In fact, this approach applies to a general parameter $W(d)$ with bounds that are minimum and maximums of linear combinations of $p(y,d|z)$’s, such as parameters that have the following form:
@@
for some linear functional $f$, where $q_{d}(y)\equiv P(Y(d)=y)$.
@@
This framework can be further generalized to incorporate treatment choices adaptive to covariates or past outcomes as in Han, (2024), analogous to Section 4.3. Sometimes, this generalization prevents the researcher from deriving analytical bounds, in which case the linear programming approach can be used.

### v2

@@ 4.1 Bounds Derived from Linear Programming
To incorporate many complicated settings, suppose that $W(d)=A_{d}q$ and $p=Bq$ for some known row vector $A_{d}$ and matrix $B$, an unknown vector $q$ in a simplex $\mathcal{Q}$, and a vector $p$ that is estimable from data. Typically $q$ is a vector of probabilities of a latent variable that governs the DGP; see Balke and Pearl, (1997, 2011), Han, (2024) and Han and Yang, (2024). The linearity in this assumption is usually implied by the nature of a particular problem (e.g., discreteness). Then we have
@@
An example of this setting can be found in Han and Yang, (2024). Let $(Y,D,Z)$ be a vector of a binary outcome, treatment and instrument and let $p$ be a vector with entries $p(y,d|z)\equiv\mathbb{P}(Y=y,D=d|Z=z)$ across $(y,d,z)\in\{0,1\}^{3}$. [^5] See Han and Yang, (2024) for the use of linear programming with continuous $Y$. Suppose $W(d)=\mathbb{E}[Y(d)]$ for $d\in\{0,1\}$. Then, we can define the response type $\varepsilon\equiv(Y(1),Y(0),D(1),D(0))$ with a realized value $e\equiv(y(1),y(0),d(1),d(0))$, where $Y(d)$ denotes the potential outcome under treatment $d$ and $D(z)$ denotes the potential treatment under instrument value $z$. Let $q(e)\equiv\mathbb{P}(\varepsilon=e)$ be the latent distribution. Then
$W(d)$ $=\mathbb{P}[Y(d)=1]=\sum_{e:y(d)=1}q(e)\equiv A_{d}q,$
@@
where $\tilde{B}\equiv\left[\begin{array}[]{c}B\\ \boldsymbol{1}^{\prime}\end{array}\right]$ is a $(d_{p}+1)\times d_{q}$ matrix with $\boldsymbol{1}$ being a $d_{q}\times 1$ vector of ones, and $\tilde{p}\equiv\left[\begin{array}[]{c}p\\ 1\end{array}\right]$ is a $(d_{p}+1)\times 1$ vector. By using a vertex enumeration algorithm (e.g., Avis and Fukuda, (1991)), one can find all (or a relevant subset) of vertices of the polyhedra $\{\lambda:\tilde{B}^{\prime}\lambda\geq-A_{d}^{\prime}\}$ and $\{\lambda:\tilde{B}^{\prime}\lambda\geq A_{d}^{\prime}\}$. Let $\Lambda_{L,d}\equiv\{\lambda_{1},...,\lambda_{J_{L,d}}\}$ and $\Lambda_{U,d}\equiv\{\lambda_{1},...,\lambda_{J_{U,d}}\}$ be the sets that collect such vertices, respectively. Then, it is easy to see that $L(d)=\max_{\lambda\in\Lambda_{L,d}}-\tilde{p}^{\prime}\lambda$ and $U(d)=\min_{\lambda\in\Lambda_{U,d}}\tilde{p}^{\prime}\lambda$, and thus Assumption 3.1 holds.
@@
where $\Delta_{\tilde{d},{d}}\equiv A_{\tilde{d}}-A_{{d}}$. Analogous to the vertex enumeration argument above, let $\Lambda_{L,\tilde{d},{d}}\equiv\{\lambda_{1},...,\lambda_{J_{L,\tilde{d},{d}}}\}$ and $\Lambda_{U,\tilde{d},{d}}\equiv\{\lambda_{1},...,\lambda_{J_{U,\tilde{d},{d}}}\}$ be the sets that collect all (or a relevant subset) of vertices of the polyhedra $\{\lambda:\tilde{B}^{\prime}\lambda\geq-\Delta_{\tilde{d},{d}}^{\prime}\}$ and $\{\lambda:\tilde{B}^{\prime}\lambda\geq\Delta_{\tilde{d},{d}}^{\prime}\}$, respectively. Then, $L(\tilde{d},{d})=\max_{\lambda\in\Lambda_{L,\tilde{d},{d}}}-\tilde{p}^{\prime}\lambda$ and $U(\tilde{d},{d})=\min_{\lambda\in\Lambda_{U,\tilde{d},{d}}}\tilde{p}^{\prime}\lambda$. Let $\widehat{\mathcal{D}}=\{d:\widehat{L}(\tilde{d},{d})\leq 0,\forall\tilde{d}\neq d\}$, where $\widehat{L}(\tilde{d},{d})$ is the sample counterpart of $L(\tilde{d},{d})$ with $\widehat{\tilde{p}}\equiv\left[\begin{array}[]{c}\hat{p}\\ 1\end{array}\right]$ replacing $\tilde{p}\equiv\left[\begin{array}[]{c}p\\ 1\end{array}\right]$. Partition $\lambda$ as $\lambda=(\lambda^{1\prime},\lambda^{0})^{\prime}$ where $\lambda^{0}$ is the last element of $\lambda$. Note that $d\in\widehat{\mathcal{D}}$ if and only if
@@ 4.2 Empirical Welfare Maximization with Observational Data
Consider allocating a binary treatment based on observed covariates $X\in\mathcal{X}$. A treatment allocation rule can be defined as a function $\delta:\mathcal{X}\rightarrow\{0,1\}$ in a class of rules $\mathcal{D}$. Consider the utilitarian welfare of deploying $\delta$ relative to treating no one. The optimal allocation $\delta^{*}$ satisfies
@@
Note that $\mathbb{E}[Y(\delta(X))-Y(0)]=E\left[\delta(X)\Delta(X)\right]$, where $\Delta(X)\equiv\mathbb{E}[Y(1)-Y(0)|X]$. This problem is considered in Kitagawa and Tetenov, (2018) and Athey and Wager, (2021), among others. When only observational data for $(Y,D,X)$ are available with $D$ being endogenous, $W(\delta)$ is only partially identified unless strong treatment effect homogeneity is assumed. This problem has been studied in Kallus and Zhou, (2021); Pu and Zhang, (2021); D’Adamo, (2021); Byambadalai, (2022), among others. Using instrumental variables, one can consider bounds on the conditional ATE based on conditional versions of the bounds considered in Sections A.1 and 4.1 (i.e., Manski’s bounds and bounds produced by linear programming).
@@
$L(\delta)$ $\equiv\mathbb{E}[\delta(X)L(X)],\qquad U(\delta)\equiv\mathbb{E}[\delta(X)U(X)].$
Similarly, bounds on $W(\tilde{\delta})-W({\delta})=\mathbb{E}[(\tilde{\delta}(X)-{\delta}(X))\Delta(X)]$ can be characterized as
$L(\tilde{\delta},{\delta})$ $\equiv\mathbb{E}[(\tilde{\delta}(X)-{\delta}(X))L(X)],\quad U(\tilde{\delta},{\delta})\equiv\mathbb{E}[(\tilde{\delta}(X)-{\delta}(X))U(X)].$ (4.7)
@@
Suppose $\mathcal{X}$ is finite and $\mathcal{X}=\{x_{1},...,x_{K}\}$ where $x_{k}$ can be a vector and $K$ can potentially be large. For simplicity of exposition, suppose $\mathcal{X}=\{0,1,2\}$. Then $\mathcal{D}=\{\delta_{1},...,\delta_{8}\}$ where each $\delta_{j}$ corresponds to a mapping type from $\{0,1,2\}$ to $\{0,1\}$. To verify Assumptions 3.1 and 3.2, we proceed as follows. For given $x\in\mathcal{X}$, by arguments analogous to those in Section 4.1 (and Section A.1), bounds $L_{x}$ and $U_{x}$ on $\Delta(x)$ satisfy, for some scalars ~$\tilde{\ell}_{j}$ and $\tilde{u}_{j}$ and row vectors $\ell_{j}$ and $u_{j}$,
$L_{x}$ $=\max_{j\in\{1,...,J_{L}\}}\{\tilde{\ell}_{j}+\ell_{j}p_{x}\},\qquad U_{x}=\min_{j\in\{1,...,J_{U}\}}\{\tilde{u}_{j}+u_{j}p_{x}\},$
where $p(x)$ is the vector of $p(y,d|z,x)$’s across $(y,d,z)$ fixing $x$. Then, by Jensen’s inequality, for each $\delta\in\mathcal{D}$,
$L(\delta)$ $\geq\tilde{L}(\delta)\equiv\max_{j\in\{1,...,J_{L}\}}\{\tilde{\ell}_{j}\mathbb{E}[\delta(X)]+\ell_{j}\mathbb{E}[\delta(X)p(X)]\},$
$U(\delta)$ $\leq\tilde{U}(\delta)\equiv\min_{j\in\{1,...,J_{U}\}}\{\tilde{u}_{j}\mathbb{E}[\delta(X)]+u_{j}\mathbb{E}[\delta(X)p(X)]\}.$
Note that $\tilde{L}(\delta)$ and $\tilde{U}(\delta)$ are non-sharp bounds; for calculation of sharp bounds, see Section A.2. We can verify Assumption 3.1 with $\tilde{L}(\delta)$ and $\tilde{U}(\delta)$ by defining
$p$ $=\left(\begin{array}[]{c}\mathbb{E}[\delta_{1}(X)]\\ \mathbb{E}[\delta_{1}(X)p(X)]\\ \vdots\\ \mathbb{E}[\delta_{8}(X)]\\ \mathbb{E}[\delta_{8}(X)p(X)]\end{array}\right)$
and, for $\delta=\delta_{1}$ as an example, by using $\ell_{\delta_{1},j}=(\begin{array}[]{ccccc}\tilde{\ell}_{j}&\ell_{j}&0&\cdots&0\end{array})$. Similarly, we can verify Assumptions 3.2 and 3.4 by estimating $p(X)$ and $\mathbb{E}[\delta(X)]$ with sample means and ${E}[\delta(X){p}(X)]$ with $\frac{1}{n}\sum_{i}^{n}\delta(X_{i})\hat{p}(X_{i}).$ If the data $\{Y_{i},D_{i},Z_{i},X_{i}\}$${}_{i=1}^{n}$ form a random sample, and $\widehat{\mathcal{D}}=\{\delta\in\mathcal{D}:\widehat{L}(\tilde{\delta},\delta)\leq 0\,\forall\tilde{\delta}\neq\delta\}$ for $\widehat{L}(\tilde{\delta},\delta)$ defined the same as ${L}(\tilde{\delta},\delta)$ in (4.7) after substituting $\hat{p}$ for $p$, Assumptions 3.2 and 3.4 hold.
@@ 4.3 Bounds for Dynamic Treatment Effects
Let $T=2$. Then $Y\equiv(Y_{1},Y_{2})$ and $D\equiv(D_{1},D_{2})$. For given welfare $W(d)$ with $d\equiv(d_{1},d_{2})$, we are interested in the optimal policy $d^{*}$ that satisfies
@@
where $\mathcal{D}\equiv\{(1,1),(1,0),(0,1),(0,0)\}$. The sign of the welfare difference, $W(d)-W(\tilde{d})$ for $d,\tilde{d}\in\mathcal{D}$, is useful for establishing the ordering of $W(d)$ with respect to $d$ and thus to identifying $d^{*}$. However, without additional identifying assumptions, we can only establish a partial ordering of $W(d)$ based on the bounds on the welfare difference (Han,, 2024). This will produce the identified set $\mathcal{D}^{*}$ for $d^{*}$.
An example of the welfare is $W(d)\equiv\mathbb{E}[Y_{2}(d)]$, namely, the average potential terminal outcome. The bounds on welfare $W(d)$ are
$L(d)$ $\equiv\max_{z}L(d;z),\qquad U(d)\equiv\min_{z}U(d;z),$ (4.8)
@@
$L(d;z)$ $\equiv\mathbb{P}(Y_{2}=1,D=d|Z=z),$
$U(d;z)$ $\equiv\mathbb{P}(Y_{2}=1,D=d|Z=z)+\sum_{d^{\prime}\neq d}\mathbb{P}(D=d^{\prime}|Z=z),$
@@
$W(\tilde{d})-W({d})=\mathbb{E}[Y_{2}(\tilde{d})-Y_{2}({d})]$ $=\mathbb{P}(Y_{2}(\tilde{d})=1)-\mathbb{P}(Y_{2}({d})=1).$
@@
$L(\tilde{d},{d})$ $\equiv L(\tilde{d})-U({d}),\qquad U(\tilde{d},{d})\equiv U(\tilde{d})-L({d}).$ (4.9)
Another example of welfare is the joint distribution $W(d)\equiv\mathbb{P}(Y(d)=(1,1))$ where $Y(d)\equiv(Y_{1}(d_{1}),Y_{2}(d))$. The bounds on $W(d)$ in this case are $L(d)\equiv\max_{z}L(d;z)$ and $U(d)\equiv\min_{z}U(d;z)$ where
$L(d;z)$ $\equiv\mathbb{P}(Y_{2}=1,D=d|Z=z),$
$U(d;z)$ $\equiv\mathbb{P}(Y_{2}=1,D=d|Z=z)+\mathbb{P}(Y_{1}=1,D_{1}=d_{1},D_{2}=d_{2}^{\prime}|Z=z)$
$\quad+\mathbb{P}(D_{1}=d_{1}^{\prime},D_{2}=d_{2}|Z=z)+\mathbb{P}(D_{1}=d_{1}^{\prime},D_{2}=d_{2}^{\prime}|Z=z),$
with $d^{\prime}_{1}\neq d_{1}$ and $d^{\prime}_{2}\neq d_{2}$. Consider the effect of treatment on the joint distribution, for example,
$W(1,1)-W(1,0)$ $=\mathbb{P}(Y(1,1)=(1,1))-\mathbb{P}(Y(1,0)=(1,1)).$
@@
$L(\tilde{d},{d})$ $\equiv\max_{z}\mathbb{P}(Y_{2}=1,D=(1,1)|Z=z)$
$\quad-\min_{z}\left\{\mathbb{P}(Y_{2}=1,D=(1,0)|Z=z)+\mathbb{P}(Y_{1}=1,D=(1,1)|Z=z)\right.$
$\left.\quad+\mathbb{P}(D=(0,1)|Z=z)+\mathbb{P}(D=(0,0)|Z=z)\right\},$
$U(\tilde{d},{d})$ $\equiv\min_{z}\left\{\mathbb{P}(Y_{2}=1,D=(1,1)|Z=z)+\mathbb{P}(Y_{1}=1,D=(1,0)|Z=z)\right.$
$\left.\quad+\mathbb{P}(D=(0,1)|Z=z)+\mathbb{P}(D=(0,0)|Z=z)\right\}$
$\quad-\max_{z}\mathbb{P}(Y_{2}=1,D=(1,0)|Z=z).$
@@
$\mathcal{D}^{*}$ =$=\{d:\nexists\tilde{d}\neq d\text{ such that }L(\tilde{d},d)>0\}=\{d:L(\tilde{d},d)\leq 0,\forall\tilde{d}\neq d\}.$
These examples are special cases of the model in Han, (2024). [^6] See also Han and Lee, (2024) for other examples of dynamic causal parameters that can be used to define optimal treatments.
In both cases, it is easy to see that $L(d)$ and $U(d)$ satisfy Assumption 3.1 with $p$ being the vector of probabilities $p(y,d|z)\equiv\mathbb{P}(Y=y,D=d|Z=z)$. To verify Assumption 3.2, let $\widehat{L}(\tilde{d},d)$ be the estimator of $L(\tilde{d},d)$ where the sample frequency replaces the population probability. Then $\widehat{\mathcal{D}}=\{d:\widehat{L}(\tilde{d},d)\leq 0,\forall\tilde{d}\neq d\}$.
@@
$\{d\in\widehat{\mathcal{D}},\widehat{z}(d)=z^{*}\}=\{\widehat{L}(\tilde{d},d)\leq 0\,\forall\tilde{d}\neq d,\widehat{z}(d)=z^{*}\}$ =$=\{A^{L}\hat{p}\leq 0\}$
@@
and, for example, $\widehat{z}(d)=1$ if and only if $\widehat{L}(d;0)-\widehat{L}(d;1)\leq 0$. A similar formulation follows for the second example. In fact, this approach applies to a general parameter $W(d)$ with bounds that are minimum and maximums of linear combinations of $p(y,d|z)$’s, such as parameters that have the following form:
@@
for some linear functional $f$, where $q_{d}(y)\equiv\mathbb{P}(Y(d)=y)$.
@@
This framework can be further generalized to incorporate treatment choices adaptive to covariates or past outcomes as in Han, (2024), analogous to Section 4.2. This generalization is considered in our empirical application in Section 8. Sometimes this generalization prevents the researcher from deriving analytical bounds, in which case the linear programming approach can be used.